# Rates of convergence in long time asymptotics of an alignment model with symmetry breaking

### v2

@@ Abstract.
$\frac{\partial f}{\partial t}=D\,\Delta f+\nabla\cdot\big{[}\left(\nabla\psi_{\alpha}(v)+v-u_{f}\right)f\big{]}\quad\text{with}\quad\psi_{\alpha}(v)=\frac{\alpha}{4}\,|v|^{4}-\frac{\alpha}{2}\,|v|^{2}.$ (1)
Here $t\geq 0$ denotes the time variable, $v\in\mathbb{R}^{d}$ is the velocity variable, $D>0$ is the intensity of the noise,  $\nabla\cdot$ and  $\Delta$ respectively denote the divergence and the Laplacian with respect to  $v$, while  $\alpha>0$ is a parameter of the external potential and  $u_{f}$ denotes the average velocity
$u_{f}:=\frac{\int_{\mathbb{R}^{d}}v\,f(t,v)\,dv}{\int_{\mathbb{R}^{d}}f(t,v)\,dv}.$
In contrast with the classical Vicsek model, we do not require that  $|v|$ is constant. For any  $u\in\mathbb{R}^{d}$, we define a Gibbs state as
$G_{u}(v):=\mathcal{Z}(u)^{-1}{e^{-\frac{1}{D}\left(\frac{1}{2}|v-u|^{2}+\psi_{\alpha}(v)\right)}}$ (2)
where  $\mathcal{Z}(u)$ is a renormalization constant. A stationary state is a Gibbs state with a specific value of the speed  $|u|$. In [2] it was proved that a phase transition occurs and the following refined result is taken from [11].
@@
There exists  $D_{*}>0$ such that:
@@
  • If $D<D_{*}$, then there exists $r(D)>0$ such that $G_{u}$ is a stationary state for any $u\in\mathcal{S}\cup\{0\}$ where
@@
    and there are no other stationary states. Moreover, $G_{0}$ is unstable and the set of polarized states $\{G_{u}\}_{u\in\mathcal{S}}$ is stable.
@@
$\mathcal{F}[f]=D\int_{\mathbb{R}^{d}}f\log f\,dv+\int_{\mathbb{R}^{d}}f\,\psi_{\alpha}\,dv+\frac{1}{2}\int_{\mathbb{R}^{d}}|v-u_{f}|^{2}\,f\,dv.$ (4)
@@
datum  $f_{\mathrm{\mathrm{ini}}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that  $\mathcal{F}[f_{\mathrm{\mathrm{ini}}}]<\mathcal{F}[G_{0}]$ and  $\|f_{\mathrm{ini}}\|_{L^{1}(\mathbb{R}^{d})}=1$. Then the average speed $u_{f}$ converges as $t\rightarrow\infty$ to a unique limit $u_{\infty}\in\mathcal{S}$ and for any $\beta>0$,
$\lim_{t\rightarrow\infty}t^{\beta}\left(\mathcal{H}\big{[}f(t,\cdot)\mid G_{u_{\infty}}\big{]}+|u_{f}(t)-u_{\infty}|^{2}\right)=0$
It was known from [11] that $\lim_{t\rightarrow\infty}\inf_{u\in\mathcal{S}}{\mathcal{H}\big{[}f(t,.)|G_{u}\big{]}}=0$ but the existence of a unique limiting Gibbs state and the convergence of $u_{f}(t)$ as $t\rightarrow\infty$ are new, thus answering a long standing open question. Under additional conditions on the initial datum, we obtain an exponential rate of convergence.
@@
There is a positive constant $\beta$ depending only on $\alpha$ and $D$ such that under the assumptions of Theorem˜1.1, if $f_{\mathrm{ini}}\in\bigcap_{u\in\mathcal{S}}L^{2}(\mathbb{R}^{d},G_{u}^{-1}dv)$, then
$\mathcal{H}\big{[}f(t,.)|G_{u_{\infty}}\big{]}+|u_{f}-u_{\infty}|^{2}=O(e^{-\beta t})\text{ as }t\rightarrow+\infty.$
The additional assumption is slightly stronger than asking $f_{\mathrm{ini}}$ to belong to $L^{2}(\mathbb{R}^{d},G_{u}^{-1}dv)$ for only one  $u$. Indeed, the function $v\mapsto\sqrt{G_{u}(v)/\left(1+|v|^{d+1}\right)}$ belongs to $L^{2}(\mathbb{R}^{d},G_{u}^{-1}\,\mathrm{d}v)$ but not to $L^{2}(\mathbb{R}^{d},G_{w}^{-1}\,\mathrm{d}v)$ if $w\neq u$. The proofs of Theorem˜1.1 and Theorem˜1.2 are based on a detailed analysis of the free energy. The exponential convergence is obtained by considering the projection of $u_{f}(t)$ onto $\mathcal{S}$.
@@ 2.1. Existence and uniqueness
We provide here some already known results and references. Since Equation˜1 has a gradient flow structure, a natural approach is to consider a discrete gradient descent scheme in the Wasserstein space, as introduced by Jordan, Kinderlehrer, and Otto in [10]. The drawback of this method is that it does not give uniqueness since the free energy is not convex. The existence is detailed in Chapter 8 of [15]. The proof only requires an initial datum in $L^{1}(\mathbb{R}^{d},\,dx)$ with finite entropy and second moment. In [3], Bolley, Cañizo and Carrillo proved the existence and uniqueness for a large class of kinetic PDE’s which covers (1). However, their method requires some additional integrability conditions on the initial datum, see [3] Theorem 1.2. In [3, 15], the solutions are measure-valued in the sense of distributions at $t=0$, and admit a density with respect to the Lebesgue measure for any $t>0$. Here, we assume that $f_{\mathrm{ini}}\in L^{1}(\mathbb{R}^{d})$ is nonnegative with unit mass and has finite free energy and second moment so that a solution exists. We use the same notation as in [11].
@@
$\mathcal{I}[f(t,.)]=D^{2}\int_{\mathbb{R}^{d}}\left|\nabla\log\left(\frac{f}{G_{u_{f}}}\right)\right|^{2}f\,dv$
@@
implies $f=G_{u_{f}}$, that is, $f=G_{u}$ with the condition of compatibility
$\mathbf{H}(u):=u-u_{G_{u}}=0.$ (6)
Conversely, by substituting $G_{u}$ into Equation˜1, one can check that $G_{u}$ is a stationary state
if $\mathbf{H}(u)=0$. We can write for $u\neq 0$
$\mathbf{H}(u)=-\frac{H(|u|)}{K(|u|)}\frac{u}{|u|}$ (7)
where $H$ and $K$ are two real functions defined by
${H}(R)=\int_{\mathbb{R}^{d}}{(v_{1}-R)\,e^{-\frac{1}{D}\left(\frac{1-\alpha}{2}|v|^{2}+\frac{\alpha}{4}|v|^{4}-v_{1}R\right)}}\,dv$ (8)
and
@@
Let $f$ be the solution to Equation˜1 with initial datum  $f_{\mathrm{ini}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{ini}}]$ is finite and $\|f_{\mathrm{ini}}\|_{L^{1}(\mathbb{R}^{d})}=1$ . Then for all $k\in\mathbb{N}$,
@@
Moreover, if $f$ solves (1), then $\mathcal{F}[f(t,.)]\leq\mathcal{F}[f_{\mathrm{ini}}]$ for all $t\geq 0$ and, as a consequence,
@@
$\frac{\mathrm{d}}{\mathrm{d}t}M_{k}$ $=Dk\,(d+k-2)\,M_{k-2}-k\,M_{k}+k\,\left(\int_{\mathbb{R}^{d}}|v|^{k-2}vf\,dv\right)\cdot u_{f}+\alpha\,k\int_{\mathbb{R}^{d}}(|v|^{k}-|v|^{k+2})f\,dv.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}M_{k}\leq D\,(d+k-2)M_{k-2}-k\,M_{k}+k\left(\int_{\mathbb{R}^{d}}|v|^{k-2}vf\,dv\right)\cdot u_{f}+\alpha\,k.$
@@ 2.5. Properties of the free energy
$\mathcal{F}_{*}:=\inf_{f\in L^{1}(\mathbb{R}^{d},(1+|v|^{4})\,\mathrm{d}v)}\mathcal{F}[f]=\begin{cases}\mathcal{F}[G_{0}]&\text{if }D\geq D_{*},\\ \mathcal{F}[G_{u}]\text{ with }|u|=r(D)&\text{if }D<D_{*}.\end{cases}$
The value of $\mathcal{F}[G_{u}]$ does not depend on $u\in\mathcal{S}$ defined by (3) since the free energy is invariant under rotations. Let us state here some useful identities. By definition, the normalized constant of a Gibbs state is
$\mathcal{Z}(u):=\int_{\mathbb{R}^{d}}e^{-\frac{1}{D}\left(\frac{1}{2}|v-u|^{2}+\psi_{\alpha}(v)\right)}\,dv.$ (11)
With a direct computation, we obtain for all $u\in\mathbb{R}^{d}$,
$\mathcal{F}[f]=\,D\,\mathcal{H}[f|G_{u}]-\frac{1}{2}\,|u-u_{f}|^{2}-D\log(\mathcal{Z}(u))\,.$ (12)
First using (12) with $u=u_{f}$ and then with $f=G_{u}$, we obtain thanks to (6) that
$\mathcal{F}[f]-\mathcal{F}_{*}=D\,\mathcal{H}[f|G_{u_{f}}]+\mathcal{V}(u_{f})-\mathcal{V}_{*}$ (13)
@@
$\mathcal{V}(u):=-\,D\log(\mathcal{Z}(u))$ (14)
is a radial function and $\mathcal{V}_{*}$ denotes the value of $\mathcal{V}$ on $\mathcal{S}$. This expression of the free energy is crucial for proving Proposition˜3.1. Taking $u\in\mathcal{S}$ in Equation˜12, we obtain
$\mathcal{F}[f]-\mathcal{F}_{*}=D\,\mathcal{H}[f|G_{u}]-\frac{1}{2}|u_{f}-u|^{2}.$ (15)
@@ 2.6. Log-Sobolev and Poincaré inequality
Let us state two inequalities which generalize the classical Gaussian Poincaré and Log-Sobolev inequalities. Since  $\phi_{u}(v):=\frac{1}{2}|v-u|^{2}+\frac{\alpha}{4}|v|^{4}-\frac{\alpha}{2}|v|^{2}$ is a bounded perturbation of a uniformly convex potential, it follows from the carré du champ method of  [1] and the Holley-Strook perturbation lemma  [9] that the Log-Sobolev and Poincaré inequalities apply with the measure $d\mu=G_{u}\,dv$.
Poincaré inequality: For all  $u\in\mathbb{R}^{d}$ there exists  $\Lambda_{u}>0$ such that for all  $g\in H^{1}(\mathbb{R}^{d},G_{u}\,dv)$ satisfying $\int_{\mathbb{R}^{d}}g\,G_{u}\,dv=0$, there holds
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{u}\,dv\geq\Lambda_{u}\int_{\mathbb{R}^{d}}g^{2}G_{u}\,dv.$ (P)
Since we will only use the Poincaré inequality with  $u\in\mathcal{S}$ we will denote by  $\Lambda(D)$ the common value of  $\Lambda_{u}$ for all  $u\in\mathcal{S}$.
Log-Sobolev inequality: For all  $u\in\mathbb{R}^{d}$, there exists  $\mathcal{K}_{u}>0$ such that for all nonnegative function of unit mass  $f$ that satisfy $\sqrt{fG_{u}^{-1}}\in H^{1}(\mathbb{R}^{d},G_{u}dv)$ there holds
$\int_{\mathbb{R}^{d}}\left|\nabla\log\left(\frac{f}{G_{u}}\right)\right|^{2}f\,dv\geq\mathcal{K}_{u}\int_{\mathbb{R}^{d}}\log\left(\frac{f}{G_{u}}\right)f\,dv.$ (LSI)
In the following, we will use the Log-Sobolev inequality with  $u=u_{f}$ in a neighborhood of  $\mathcal{S}$.
Let  $\mathcal{K}_{\epsilon}=\inf_{u\in\mathbb{R}^{d}\,:\,\mathrm{dist}(u,\mathcal{S})\leq\epsilon}\mathcal{K}_{u}>0$ where $\mathrm{dist}(u,\mathcal{S})=\min_{v\in\mathcal{S}}|u-v|$.
@@
Let $D>0$. Let $f$ be the solution to Equation˜1 with initial datum  $f_{\mathrm{ini}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{ini}}]<+\infty$ and $\|f_{\mathrm{ini}}\|_{L^{1}(\mathbb{R}^{d})}=1$.
@@
    $\lim_{t\rightarrow{\infty}}u_{f}(t)=0\,.$
  • If $D<D_{*}$ and $\mathcal{F}[f_{\mathrm{ini}}]<\mathcal{F}[G_{0}]$ we have
    $\lim_{t\rightarrow{\infty}}\mathrm{dist}(u_{f}(t),\mathcal{S})=0\,.$
@@
$\lim_{t\rightarrow\infty}|u_{f}(t)-u_{G_{u_{f}}}(t)|=0.$
Let $\gamma(t)=|u_{f}(t)-u_{G_{u_{f}}}(t)|^{\beta}$ with $\beta>0$.
$\int_{0}^{\infty}\gamma(t)dt$ $=\int_{0}^{\infty}\left|\int_{\mathbb{R}^{d}}(vf-v\,G_{u_{f}})\,dv\right|^{\beta}dt$
$\leq\int_{0}^{\infty}\left[\left(\int_{\mathbb{R}^{d}}\,|f-G_{u_{f}}|\,dv\right)^{\frac{3}{4}}\left(\int_{\mathbb{R}^{d}}|v|^{4}\,|f+G_{u_{f}}|\,dv\right)^{\frac{1}{4}}\right]^{\beta}dt$
$\leq C\int_{0}^{\infty}\left(\int_{\mathbb{R}^{d}}\,|f-G_{u_{f}}|\,dv\right)^{\frac{3\beta}{4}}dt$
@@
$\int_{0}^{\infty}\gamma(t)dt\leq C\int_{0}^{\infty}\left(\int_{\mathbb{R}^{d}}|f-G_{u_{f}}|dv\right)^{2}.$ (16)
@@
$\frac{1}{2}\left\|f-G_{u_{f}}\right\|^{2}_{L^{1}(\mathbb{R}^{d})}\leq\mathcal{H}[f|G_{u_{f}}]$ (C-K)

### v3

@@ Abstract.
$\frac{\partial f}{\partial t}=D\,\Delta f+\nabla\cdot\big[\left(\nabla\psi_{\alpha}(v)+v-\mathbf{u}_{f}\right)f\big]\quad\text{with}\quad\psi_{\alpha}(v)=\frac{\alpha}{4}\,|v|^{4}-\frac{\alpha}{2}\,|v|^{2}.$ (1)
Here $t\geq 0$ denotes the time variable, $v\in\mathbb{R}^{d}$ is the velocity variable, $D>0$ is the intensity of the noise, $\nabla\cdot$ and $\Delta$ respectively denote the divergence and the Laplacian with respect to $v$, while $\alpha>0$ is a parameter of the external potential and $\mathbf{u}_{f}$ denotes the average velocity
$\mathbf{u}_{f}:=\frac{\int_{\mathbb{R}^{d}}v\,f(t,v)\,dv}{\int_{\mathbb{R}^{d}}f(t,v)\,dv}.$
In contrast with the classical Vicsek model, we do not require that $|v|$ is constant. For any $\mathbf{u}\in\mathbb{R}^{d}$, we define a Gibbs state as
$G_{\mathbf{u}}(v):=\mathcal{Z}(\mathbf{u})^{-1}{e^{-\frac{1}{D}\left(\frac{1}{2}|v-\mathbf{u}|^{2}+\psi_{\alpha}(v)\right)}}$ (2)
where $\mathcal{Z}(\mathbf{u})$ is a renormalization constant. A stationary state is a Gibbs state with a specific value of the speed $|\mathbf{u}|$. In [2] it was proved that a phase transition occurs and the following refined result is taken from [11].
@@
Let $D_{*}>0$ be the unique positive root of
$h(D):=\int_{0}^{\infty}\left(s^{d+1}-s^{d+3}\right)e^{-\frac{\psi_{\alpha}(v)}{D}-\frac{|v|^{2}}{2D}}\,dv.$
For $D<D_{*},$ let $r(D)>0$ be the unique positive root of
${H}(R):=\int_{\mathbb{R}^{d}}{(v_{1}-R)\,e^{-\frac{1}{D}\left(\psi_{\alpha}(v)+\frac{|v|^{2}}{2}-v_{1}R\right)}}\,dv.$ (3)
@@
  • If $D<D_{*}$, then $G_{\mathbf{u}}$ is a stationary state for any $u\in\mathcal{S}\cup\{0\}$ where
@@
    and there are no other stationary states. Moreover, $G_{0}$ is unstable and the set of polarized states $\{G_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{S}}$ is stable.
@@
$\mathcal{F}[f]=D\int_{\mathbb{R}^{d}}f\log f\,dv+\int_{\mathbb{R}^{d}}f\,\psi_{\alpha}\,dv+\frac{1}{2}\int_{\mathbb{R}^{d}}|v-\mathbf{u}_{f}|^{2}\,f\,dv.$ (5)
@@
datum $f_{\mathrm{\mathrm{in}}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{\mathrm{in}}}]<\mathcal{F}[G_{0}]$ and $\|f_{\mathrm{in}}\|_{L^{1}(\mathbb{R}^{d})}=1$. Then the average speed $\mathbf{u}_{f}$ converges as $t\rightarrow\infty$ to a unique limit $\mathbf{u}_{\infty}\in\mathcal{S}$ and for any $\beta>0$,
$\lim_{t\rightarrow\infty}t^{\beta}\left(\mathcal{H}\big[f(t,\cdot)\mid G_{\mathbf{u}_{\infty}}\big]+|\mathbf{u}_{f}(t)-\mathbf{u}_{\infty}|^{2}\right)=0$
It was known from [11] that $\lim_{t\rightarrow\infty}\inf_{\mathbf{u}\in\mathcal{S}}{\mathcal{H}\big[f(t,.)|G_{\mathbf{u}}\big]}=0$ but the existence of a unique limiting Gibbs state and the convergence of $\mathbf{u}_{f}(t)$ as $t\rightarrow\infty$ are new, thus answering a long standing open question. Under additional conditions on the initial datum, we obtain an exponential rate of convergence.
@@
There is a positive constant $\beta$ depending only on $\alpha$ and $D$ such that under the assumptions of Theorem˜1.1, if $f_{\mathrm{in}}\in\bigcap_{u\in\mathcal{S}}L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}^{-1}dv)$, then
$\mathcal{H}\big[f(t,.)|G_{\mathbf{u}_{\infty}}\big]+|\mathbf{u}_{f}-\mathbf{u}_{\infty}|^{2}=O(e^{-\beta t})\text{ as }t\rightarrow+\infty.$
The additional assumption is slightly stronger than asking $f_{\mathrm{in}}$ to belong to $L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}^{-1}dv)$ for only one $\mathbf{u}$. Indeed, the function $v\mapsto\sqrt{G_{\mathbf{u}}(v)/\left(1+|v|^{d+1}\right)}$ belongs to $L^{2}(\mathbb{R}^{d},G_{\mathbf{u}}^{-1}\,\mathrm{d}v)$ but not to $L^{2}(\mathbb{R}^{d},G_{w}^{-1}\,\mathrm{d}v)$ if $w\neq\mathbf{u}$. The proofs of Theorem˜1.1 and Theorem˜1.2 are based on a detailed analysis of the free energy. The exponential convergence is obtained by considering the projection of $\mathbf{u}_{f}(t)$ onto $\mathcal{S}$.
@@ 2.1. Existence and uniqueness
We provide here some already known results and references. Since Equation˜1 has a gradient flow structure, a natural approach is to consider a discrete gradient descent scheme in the Wasserstein space, as introduced by Jordan, Kinderlehrer, and Otto in [10]. The drawback of this method is that it does not give uniqueness since the free energy is not convex. The existence is detailed in Chapter 8 of [15]. The proof only requires an initial datum in $L^{1}(\mathbb{R}^{d},\,dx)$ with finite entropy and second moment. In [3], Bolley, Cañizo and Carrillo proved the existence and uniqueness for a large class of kinetic PDE’s which covers (1). However, their method requires some additional integrability conditions on the initial datum, see [3] Theorem 1.2. In [3, 15], the solutions are measure-valued in the sense of distributions at $t=0$, and admit a density with respect to the Lebesgue measure for any $t>0$. Here, we assume that $f_{\mathrm{in}}\in L^{1}(\mathbb{R}^{d})$ is nonnegative with unit mass and has finite free energy and second moment so that a solution exists. We use the same notation as in [11].
@@
$\mathcal{I}[f(t,.)]=D^{2}\int_{\mathbb{R}^{d}}\left|\nabla\log\left(\frac{f}{G_{\mathbf{u}_{f}}}\right)\right|^{2}f\,dv$
@@
implies $f=G_{\mathbf{u}_{f}}$, that is, $f=G_{\mathbf{u}}$ with the condition of compatibility
$\mathbf{H}(\mathbf{u}):=\mathbf{u}-\mathbf{u}_{G_{\mathbf{u}}}=0.$ (7)
Conversely, by substituting $G_{\mathbf{u}}$ into Equation˜1, one can check that $G_{\mathbf{u}}$ is a stationary state
if $\mathbf{H}(\mathbf{u})=0$. We can write for $\mathbf{u}\neq 0$
$\mathbf{H}(\mathbf{u})=-\frac{H(|\mathbf{u}|)}{K(|\mathbf{u}|)}\frac{\mathbf{u}}{|\mathbf{u}|}$ (8)
where $H$ and $K$ are two real functions defined by (3) and
@@
Let $f$ be the solution to Equation˜1 with initial datum $f_{\mathrm{in}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{in}}]$ is finite and $\|f_{\mathrm{in}}\|_{L^{1}(\mathbb{R}^{d})}=1$ . Then for all $k\in\mathbb{N}$,
@@
Moreover, if $f$ solves (1), then $\mathcal{F}[f(t,.)]\leq\mathcal{F}[f_{\mathrm{in}}]$ for all $t\geq 0$ and, as a consequence,
@@
$\frac{\mathrm{d}}{\mathrm{d}t}M_{k}$ $=Dk\,(d+k-2)\,M_{k-2}-k\,M_{k}+k\,\left(\int_{\mathbb{R}^{d}}|v|^{k-2}vf\,dv\right)\cdot\mathbf{u}_{f}+\alpha\,k\int_{\mathbb{R}^{d}}(|v|^{k}-|v|^{k+2})f\,dv.$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}M_{k}\leq D\,(d+k-2)M_{k-2}-k\,M_{k}+k\left(\int_{\mathbb{R}^{d}}|v|^{k-2}vf\,dv\right)\cdot\mathbf{u}_{f}+\alpha\,k.$
@@ 2.5. Properties of the free energy
$\mathcal{F}_{*}:=\inf_{f\in L^{1}(\mathbb{R}^{d},(1+|v|^{4})\,\mathrm{d}v)}\mathcal{F}[f]=\begin{cases}\mathcal{F}[G_{0}]&\text{if }D\geq D_{*},\\ \mathcal{F}[G_{\mathbf{u}}]\text{ with }|u|=r(D)&\text{if }D<D_{*}.\end{cases}$
The value of $\mathcal{F}[G_{\mathbf{u}}]$ does not depend on $u\in\mathcal{S}$ defined by (4) since the free energy is invariant under rotations. Let us state here some useful identities. By definition, the normalized constant of a Gibbs state is
$\mathcal{Z}(\mathbf{u}):=\int_{\mathbb{R}^{d}}e^{-\frac{1}{D}\left(\frac{1}{2}|v-\mathbf{u}|^{2}+\psi_{\alpha}(v)\right)}\,dv.$ (11)
With a direct computation, we obtain for all $\mathbf{u}\in\mathbb{R}^{d}$,
$\mathcal{F}[f]=\,D\,\mathcal{H}[f|G_{\mathbf{u}}]-\frac{1}{2}\,|\mathbf{u}-\mathbf{u}_{f}|^{2}-D\log(\mathcal{Z}(\mathbf{u}))\,.$ (12)
First using (12) with $\mathbf{u}=\mathbf{u}_{f}$ and then with $f=G_{\mathbf{u}}$, we obtain thanks to (7) that
$\mathcal{F}[f]-\mathcal{F}_{*}=D\,\mathcal{H}[f|G_{\mathbf{u}_{f}}]+\mathcal{V}(\mathbf{u}_{f})-\mathcal{V}_{*}$ (13)
@@
$\mathcal{V}(\mathbf{u}):=-\,D\log(\mathcal{Z}(\mathbf{u}))$ (14)
is a radial function and $\mathcal{V}_{*}$ denotes the value of $\mathcal{V}$ on $\mathcal{S}$. This expression of the free energy is crucial for proving Proposition˜3.1. Taking $\mathbf{u}\in\mathcal{S}$ in Equation˜12, we obtain
$\mathcal{F}[f]-\mathcal{F}_{*}=D\,\mathcal{H}[f|G_{\mathbf{u}}]-\frac{1}{2}|\mathbf{u}_{f}-\mathbf{u}|^{2}.$ (15)
@@ 2.6. Log-Sobolev and Poincaré inequality
Let us state two inequalities which generalize the classical Gaussian Poincaré and Log-Sobolev inequalities. Since $\phi_{\mathbf{u}}(v):=\frac{1}{2}|v-\mathbf{u}|^{2}+\frac{\alpha}{4}|v|^{4}-\frac{\alpha}{2}|v|^{2}$ is a bounded perturbation of a uniformly convex potential, it follows from the carré du champ method of [1] and the Holley-Strook perturbation lemma [9] that the Log-Sobolev and Poincaré inequalities apply with the measure $d\mu=G_{\mathbf{u}}\,dv$.
Poincaré inequality: For all $\mathbf{u}\in\mathbb{R}^{d}$ there exists $\Lambda_{\mathbf{u}}>0$ such that for all $g\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}\,dv)$ satisfying $\int_{\mathbb{R}^{d}}g\,G_{\mathbf{u}}\,dv=0$, there holds
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{\mathbf{u}}\,dv\geq\Lambda_{\mathbf{u}}\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}\,dv.$ (P)
Since we will only use the Poincaré inequality with $\mathbf{u}\in\mathcal{S}$ we will denote by $\Lambda(D)$ the common value of $\Lambda_{\mathbf{u}}$ for all $\mathbf{u}\in\mathcal{S}$.
Log-Sobolev inequality: For all $\mathbf{u}\in\mathbb{R}^{d}$, there exists $\mathcal{K}_{\mathbf{u}}>0$ such that for all nonnegative function of unit mass $f$ that satisfy $\sqrt{fG_{\mathbf{u}}^{-1}}\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}dv)$ there holds
$\int_{\mathbb{R}^{d}}\left|\nabla\log\left(\frac{f}{G_{\mathbf{u}}}\right)\right|^{2}f\,dv\geq\mathcal{K}_{\mathbf{u}}\int_{\mathbb{R}^{d}}\log\left(\frac{f}{G_{\mathbf{u}}}\right)f\,dv.$ (LSI)
In the following, we will use the Log-Sobolev inequality with $\mathbf{u}=\mathbf{u}_{f}$ in a neighborhood of $\mathcal{S}$.
Let $\mathcal{K}_{\epsilon}=\inf_{\mathbf{u}\in\mathbb{R}^{d}\,:\,\mathrm{dist}(\mathbf{u},\mathcal{S})\leq\epsilon}\mathcal{K}_{\mathbf{u}}>0$ where $\mathrm{dist}(\mathbf{u},\mathcal{S})=\min_{v\in\mathcal{S}}|\mathbf{u}-v|$.
@@
Let $D>0$. Let $f$ be the solution to Equation˜1 with initial datum $f_{\mathrm{in}}\in L^{1}_{+}(\mathbb{R}^{d})$ such that $\mathcal{F}[f_{\mathrm{in}}]<+\infty$ and $\|f_{\mathrm{in}}\|_{L^{1}(\mathbb{R}^{d})}=1$.
@@
    $\lim_{t\rightarrow{\infty}}\mathbf{u}_{f}(t)=0\,.$
  • If $D<D_{*}$ and $\mathcal{F}[f_{\mathrm{in}}]<\mathcal{F}[G_{0}]$ we have
    $\lim_{t\rightarrow{\infty}}\mathrm{dist}(\mathbf{u}_{f}(t),\mathcal{S})=0\,.$
@@
$\lim_{t\rightarrow\infty}|\mathbf{u}_{f}(t)-\mathbf{u}_{G_{\mathbf{u}_{f}}}(t)|=0.$
Let $\gamma(t)=|\mathbf{u}_{f}(t)-\mathbf{u}_{G_{\mathbf{u}_{f}}}(t)|^{\beta}$ with $\beta>0$.
$\int_{0}^{\infty}\gamma(t)dt$ $=\int_{0}^{\infty}\left|\int_{\mathbb{R}^{d}}(vf-v\,G_{\mathbf{u}_{f}})\,dv\right|^{\beta}dt$
$\leq\int_{0}^{\infty}\left[\left(\int_{\mathbb{R}^{d}}\,|f-G_{\mathbf{u}_{f}}|\,dv\right)^{\frac{3}{4}}\left(\int_{\mathbb{R}^{d}}|v|^{4}\,|f+G_{\mathbf{u}_{f}}|\,dv\right)^{\frac{1}{4}}\right]^{\beta}dt$
$\leq C\int_{0}^{\infty}\left(\int_{\mathbb{R}^{d}}\,|f-G_{\mathbf{u}_{f}}|\,dv\right)^{\frac{3\beta}{4}}dt$
@@
$\int_{0}^{\infty}\gamma(t)dt\leq C\int_{0}^{\infty}\left(\int_{\mathbb{R}^{d}}|f-G_{\mathbf{u}_{f}}|dv\right)^{2}.$ (16)
@@
$\frac{1}{2}\left\|f-G_{\mathbf{u}_{f}}\right\|^{2}_{L^{1}(\mathbb{R}^{d})}\leq\mathcal{H}[f|G_{\mathbf{u}_{f}}]$ (C-K)